Math 215 Section 2 Fall 95 Homework Assignments

Assignment Date Reading Problems  Due Date
8/28  1.1  1, 6, 15, 18, 24, 26  9/6
8/30  1.2  2, 6, 7, 8, 19, 20  9/6
9/1  2.1  4, 10, 16, 18, 25, 28  9/6

For extra credit (and some extra insight) try problems 30-32 of Section 1.2.